Question:

Suppose that business travelers and vacationers have the following demand for airline tickets from New York to Boston:


PriceQuantity Demanded (business travelers)Quantity Demanded (vacationers)

$150

2,100tickets

1,000tickets

200

2,000

800

250

1,900

600

300

1,800

400

(a.) As the price of tickets rises from $200 to $250, what is the price elasticity of demand for

(i) business travelers and (ii) vacationers?


1
Expert's answer
2021-10-21T12:33:17-0400

We will use the midpoint method to compute the price elasticity of demand for given categories. 

(i) business travelers

Percentage change in price

"Percentage\: change\: in\: Price\, = \frac{P2\, -\, P1}{(P2\, +\, P1)\, \div\, 2}\times 100"

  • P1 = 200
  • P2 = 250

Hence,

"Percentage\: change\: in\: Price\, = \frac{250\, -\, 200}{(250\, +\, 200)\, \div\, 2}\times 100\,=22.22\%"

Percentage change in the quantity demanded

"Percentage\: change\: in\: Quantity\, = \frac{Q2\, -\, Q1}{(Q2\, +\, Q1)\, \div\, 2}\times 100"

  • Q1 = 2000
  • Q2 = 1900

"Percentage\: change\: in\: Quantity\, = \frac{1,900 \, -\, 2,000}{(1,900 \, +\, 2000)\, \div\, 2}\times 100\,=\,-5.13\%"

Price elasticity of demand is given as:

"Price\, elasticity\, of\, demand\,= \frac{Percentage\, change\, in\, quantity\, demanded}{Percentage\, change\, in\, price}"

"Price\, elasticity\, of\, demand\,= \frac{-5.13\%\,}{22.22\%}\,=-0.23"


ii) vacationers?

Percentage change in price

From the previous computations, we know that:

Percentage change in price = 22.22%

Percentage change in the quantity demanded

Q1 = 800

Q2 = 600

"Percentage\: change\: in\: Quantity\, = \frac{600 \, -\, 800}{(600 \, +\, 800)\, \div\, 2}\times 100\,=\,-28.57\%"


"Price\, elasticity\, of\, demand\,= \frac{-28.57\%\,}{22.22\%}\,=\,-1.29"